OpenStudy (anonymous):

the sum of two consecutive odd integers is 24

OpenStudy (anonymous):

Well if they are consecutive, they come right after one another, since they are odd, you should start with an odd number less than 24 and see if the number next to it adds up to 24 when combined with it.

OpenStudy (anonymous):

(2n+1)+(2n+3) = 24

OpenStudy (jiteshmeghwal9):

Let one consecutive odd integer be 'x'.Then, sec. odd integer=(x+2) A/q, =>(x)+(x+2)=24 =>2x+2=24 Now ,can u solve it from here???

OpenStudy (jiteshmeghwal9):

@Courtnee